### v1

# Sensing Gravity with Polarized Electromagnetic Radiation

Abstract
Polarization wiggling is a direct gravitational effect exerted by a gravitational field on any electromagnetic radiation traversing it. This effect is investigated in linear gravity for spacetimes with flat or conformally flat backgrounds. First, we show how the polarization wiggle rates can be computed in the conformal frame and transformed to the physical frame. Then it is shown that polarization wiggling is not sensitive to scalar perturbations to the metric, while vector and tensor perturbations do induce polarization wiggling. This poses two natural questions: Can polarized electromagnetic radiation be used to measure vectorial and tensorial components of gravitational fields directly? And if so, how? Next, polarization wiggling is studied for an arbitrary vector perturbation to the spacetime metric. In a stationary spacetime, the polarization wiggle rate is proportional to the difference in frame dragging rates between radiation emission and measurement events. We show how this can be used to measure the angular momentum of a gravitational source if the emitter orbits the gravitational source on a known orbit. Next, the effect of a gravitational tensor mode with arbitrary polarization is studied. Finally, the polarization wiggling effect induced by a gravitational tensor mode is analyzed. Two cases are demonstrated: A spacetime with a flat Minkowski background and an expanding cosmology with a conformally flat background. In both cases, the polarization wiggling frequency equals the frequency of the gravitational tensor mode, while the other state parameters are encoded in  the polarization wiggling amplitude and phase of the polarized radiation. We show that measurements of polarization wiggling frequency, amplitude and phase of polarized radiation from different directions enables all state parameters of a gravitational tensor mode to be determined.
I Introduction
Electromagnetic radiation is our primary tool for probing the universe, and its polarization degrees of freedom represent a separate channel of information about the processes and phenomena with which the radiation interacts. While gravitational effects on electromagnetic radiation like gravitational lensing and redshift are well understood and used in astrophysical measurements, gravity’s direct effect on electromagnetic polarization is among the most elusive effects of gravity. Pending sufficiently accurate techniques for measuring gravitational effects on polarized radiation, it is important to understand them and how they can be used as separate means of probing gravitational fields, possibly beyond what is possible to do by other means.
The effect of a gravitomagnetic field on electromagnetic polarization has been studied extensively from a theoretical perspective by many authors over many years [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. The effect is known as the Gravitational Faraday Effect, alternately as the Skrotskii Effect after G. V. Skrotskii, who first described the effect [1]. From an observational perspective, however, the Gravitational Faraday Effect on a single ray of radiation provides an insufficient view of the effect of gravity on electromagnetic polarization, because it lacks unambiguous quantification [18, 23].
Defined in distinctly operational terms, polarizaton wiggling, on the other hand, provides an unambiguous quantification of the effect of gravity on the state of polarization of electromagnetic radiation [23]. It represents the effect from the perspective of a single inertial observer: Given a pointlike inertial emitter of polarized electromagnetic radation, it uniquely defines how the presence of a gravitational field along the trajectory of the radiation is manifested in observations made by this observer. Given an inertial emitter of polarized electromagnetic radiation, polarization wiggling manifests as a change in the measured state of polarization between subsequent observation events by a single inertial detector. For linearly or elliptically polarized radiation, it will be seen as a motion of the polarization axis. Polarizaton wiggling is by definition an observable gravitational effect on the state of electromagnetic polarization. It is induced by gravitational fields present along the radiation path between the radiation source and the detector.
In a recent paper, the polarization wiggling effect was studied for an arbitrary 4-dimensional spacetime and electromagnetic radiation with arbitrary state of polarization and frequency composition [23]. In the present paper, these results will be applied to study the effect of weak gravitational fields on electromagnetic polarization in more detail by using linearized gravity.
The main objective of this paper is, by the use of linear gravity in spacetimes with flat or conformally flat backgrounds, to elucidate the polarization wiggling effect through a study of the effect of scalar, vector and tensor metric perturbations on electromagnetic polarization. We will show that polarization wiggling is not sensitive to scalar metric perturbations, while vector and tensor perturbations yield independent, gauge invariant contributions to the polarization wiggle rates.
The polarization wiggling effect is reviewed in Section II. Section III is an overture to the rest of the paper in which we study how the polarization wiggle rates transform under homogeneous conformal transformations. It is shown how the  polarization wiggle rate in the physical frame is related to the polarization wiggle rate in the presumably much simpler conformal frame. This greatly simplifies later calculations and enables a study of polarization wiggling for spacetimes with conformally flat background geometries, relevant in a cosmological setting.
In Section IV, we study polarization wiggling in linear gravity. In linear gravity, the gravitational field can be described as a perturbation to a homogeneous spacetime geometry referred to as the background geometry. When linearizing a metric theory of gravity, the gravitational field can be decomposed into scalar, vector and tensor perturbations that are sourced and evolve independently. In Section V, we show how this metric decomposition leads to a gauge invariant decomposition of the polarization wiggle rate into independent components induced by scalar, vector and tensor perturbations separately.
In Section VII, we consider the effect of general vector perturbations, and show how the polarization wiggle rate is related to the gravitomagnetic field. As a particular example, we calculate the effect of a stationary, axi-symmetric gravitational field induced by a rotating body on polarized radiation emitted by an emitter orbiting the gravitational source. We show how a series of polarization wiggle rate measurements of radiation emitted at different positions along the orbit will enable the angular momentum of the gravitational source to be determined.
Finally, in Section VIII, we consider the effect of a gravitational tensor mode with an arbitrary state of polarization. Two particular examples are shown: the effect of a gravitational tensor mode propagating in a spacetime with a flat Minkowski background geometry and a gravitational tensor mode propagating in an expanding cosmology with a conformally flat background geometry. We show how radiation from multiple polarized sources can be used to determine all state parameters of the gravitational tensor mode.
Our metric convention is ($+$)(-$+$)( - + + + ), and relativistic units with $c=1$ are used. The Einstein summation convention is assumed on repeated indices.
II Polarization Wiggling
In this section, we will briefly review the polarization wiggling effect. This effect was studied in detail in a recent paper [23], and we will use the nomenclature and notation introduced there. For simplicity, we will restrict our attention to plane electromagnetic waves with frequencies above the geometric optics limit. The results generalize to composite electromagnetic radiation with arbitrary state of polarization and frequency spectrum, given that the emitted frequency spectrum has negligible power below the geometric optics limit and that the radiation is emitted with a constant frequency spectrum [23].
Polarization wiggling is a motion (wiggling) of the polarization axis of electromagnetic radiation, as observed by an inertial observer. Polarization wiggling is a pure gravitational effect that is induced by the gravitational fields present along the geodesic segment between the emitter and the detector.
First, let us assume the presence of a ray of polarized electromagnetic radiation, represented as an electromagnetic plane wave. In the geometric optics limit, it can be represented by a vector potential
$A_{\mu}=a_{\mu}e^{i\vartheta},$
where $a_{\mu}$ is a slowly varying complex amplitude and $\vartheta$ is a rapidly varying real phase. Its wave vector $p_{\mu}=\nabla_{\mu}\vartheta$ is a null vector that satisfies the null geodesic equation $\nabla_{p}p^{\mu}=0$. The magnitude of the vector potential is the amplitude $a\equiv{(A_{\mu}^{*}A^{\mu})}^{1/2}$. The polarization vector $f^{\mu}\equiv a^{\mu}/a$ is a spacelike, complex unit vector that is parallel transported along the radiation geodesic.
Next, we will assume the presence of a congruence of  timelike geodesics $u^{\mu}$. They will be referred to as the rest frame geodesics, or rest frame, for short. With this decomposition, the spatial projector is defined by
$\gamma_{\nu}^{\mu}=\delta_{\nu}^{\mu}+u^{\mu}u_{\nu}.$ (1)
This decomposition can be made for any spacetime geometry, and does not put any constraints on the spacetime geometry.
Then, turning our attention to the null geodesic congruence $p^{\mu}$, at any point along one of its geodesics, the null vector $p^{\mu}$ can be decomposed into timelike and spacelike vectors as follows:
$p^{\mu}=p(u^{\mu}+{\hat{p}}^{\mu}),$ (2)
where $p\equiv-u_{\mu}p^{\mu}$ is the photon energy measured in an inertial frame with geodesic vector $u^{\mu}$, while $\hat{p}$ is the vector
${\hat{p}}^{\mu}\equiv\frac{p^{\mu}}{p}-u^{\mu}=\gamma_{\nu}^{\mu}\frac{p^{\nu}}{p}.$ (3)
${\hat{p}}^{\mu}$ is a spacelike unit vector. It is orthogonal to $u^{\mu}$, and represents the spatial direction of propagation of the radiation. The scalar momentum can be found by contracting it with the 4-momentum: $p=p_{\mu}{\hat{p}}^{\mu}$.
The screen projector $S_{\mu\nu}$ adapted to $p^{\mu}$ is defined as
$S_{\mu\nu}=\gamma_{\mu\nu}-{\hat{p}}_{\mu}{\hat{p}}_{\nu}=g_{\mu\nu}+u_{\mu}u_{\nu}-{\hat{p}}_{\mu}{\hat{p}}_{\nu},$ (4)
while the screen rotator $\epsilon_{\mu\nu}$ is defined as
$\epsilon_{\mu\nu}\equiv\epsilon_{\alpha\beta\mu\nu}u^{\alpha}{\hat{p}}^{\beta}.$ (5)
$\epsilon_{\alpha\beta\mu\nu}$ is the Levi-Civita tensor. The screen projector $S_{\mu\nu}$ projects any 4-vector onto the plane transverse to the direction of propagation $\hat{p}$ (the screen), while the screen rotator $\epsilon_{\mu\nu}$ is a generator of rotations in the screen [23].
Expressed in terms of the Riemann tensor $R_{\gamma\rho\beta\sigma}$, the curvature twist scalar $\mathcal{Z}$ is defined as
$\mathcal{Z}\equiv u^{\gamma}{\hat{p}}^{\beta}\epsilon^{\rho\sigma}R_{\gamma\rho\beta\sigma}.$ (6)
The curvature twist $\mathcal{Z}$ can be interpreted as an angular acceleration of the polarization axis as the radiation propagates along the geodesic, the wiggle acceleration [23].
By using the relationship between the Riemann and Weyl tensors, the curvature twist can alternatively be expressed in terms of the Weyl tensor $C_{\gamma\rho\beta\sigma}$ as
$\mathcal{Z}=u^{\gamma}{\hat{p}}^{\beta}\epsilon^{\rho\sigma}C_{\gamma\rho\beta\sigma}.$ (7)
The transverse polarization vector $\epsilon^{\mu}$ is defined as the transverse projection of the polarization vector $f^{\mu}$: $\epsilon^{\mu}\equiv S_{\nu}^{\mu}f^{\nu}$. Two quantities quantifying the gravitational effect on electromagnetic polarization are the polarization axis wiggle rate $\omega$ and the mean helicity wiggle rate $\chi$. They are defined as two independent projections of the transverse time rate of change of the transverse polarization vector $\epsilon^{\mu}$, and therefore quantify the time rate of change of the observed state of polarization in an inertial frame [23]. For linearly polarized radiation or radiation emitted with constant elliptical polarization, $\omega$ equals the rotation rate of the polarization axis. $\chi$ quantifies, as the name suggests, the  mean of the rates of change of the two helicity phases of the radiation. For radiation emitted with constant elliptical polarization, $\chi$ is proportional to the rotation rate of the polarization axis. $\omega$ is independent of both radiation frequency and the emitted state of polarization, while $\chi$ depends on the degree of circular polarization, $\mathcal{V}$. $\chi$ vanishes for linearly polarized radiation, while $\omega$ is undefined for radiation with pure circular polarization. Their transport equations along the null geodesic are
$\nabla_{p}\left(\frac{\omega}{p}\right)=\mathcal{Z}$ (8)
$\nabla_{p}\left(\frac{\chi}{p}\right)=-\mathcal{V}\mathcal{Z}.$ (9)
The circular polarization degree $\mathcal{V}$ is conserved along the radiation geodesic [23]:
$\nabla_{p}\mathcal{V}=0.$ (10)
A linear combination of eqs. (8) and (9) yields
$\nabla_{p}\left(\frac{\chi+\mathcal{V}\omega}{p}\right)=0.$ (11)
The solution to eq. (8) is
$\omega(\lambda)=\frac{p(\lambda)}{p(\lambda_{*})}\omega(\lambda_{*})+p(\lambda)\text{}\operatorname*{\int}\limits_{\lambda_{*}}^{\lambda}d\lambda^{\prime}\mathcal{Z}(\lambda^{\prime}),$ (12)
where $\lambda$ denotes the affine parameter of the null geodesic of the radiation and $\lambda_{*}$ labels the emission event. Eq. (11) implies that $\chi(\lambda)$ can be expressed in terms of $\omega(\lambda)$:
$\chi(\lambda)=\frac{p(\lambda)}{p(\lambda_{*})}\left(\chi(\lambda_{*})+\mathcal{V}\omega(\lambda_{*})\right)-\mathcal{V}\omega(\lambda),$ (13)
The $p(\lambda)/p(\lambda_{*})$ factor in the first terms on the right hand side of eqs. (12) and (13) can be expressed in terms of the redshift $z(\lambda_{*})$ of the radiation source:
$\frac{p(\lambda)}{p(\lambda_{*})}={\left(1+z(\lambda_{*})\right)}^{-1}.$
For radiation emitted with constant polarization, the initial conditions are $\omega(\lambda_{*})=\chi(\lambda_{*})=0$. This summarizes previous results needed for the present paper.
III The Scale Dependency of the Polarization Wiggle Rates
Before we address the main topic of this paper, let us start with a prelude that will be used in later sections. We would like to understand how polarization wiggling is manifested in expanding cosmologies. In particular, we would like to understand how the curvature twist and the polarization wiggle rates vary with the scale factor.
In order to separate the scale dependency from the dependencies on other metric components, we will assume that the spacetime geometry of the physical spacetime $\tilde{\mathcal{M}}$ is conformally related to another geometry $\mathcal{M}$, referred to as the conformal frame geometry or simply the conformal frame. Their metrics are then related by a conformal factor $\Omega^{2}$:
$\tilde{g}=\Omega^{2}g.$
Here, $\tilde{g}$ denotes the metric of the physical spacetime geometry $\tilde{\mathcal{M}}$, while $g$ represents the metric of the conformal frame geometry $\mathcal{M}$. For example, the Minkowski geometry is the conformal frame of the conformally flat background geometries commonly used in cosmology.
We can make use of two facts to simplify the computation of polarization wiggle rates: 1) Null geodesics are invariant under conformal transformations, and 2) Maxwell’s equations in vacuum are conformally invariant. As will be shown shortly, this enables us to do all the necessary computations in the conformal frame, which is normally a much simpler geometry than the physical frame, and simply transform the results to the physical frame by using simple scale factors involving the conformal factor $\Omega$.
In the following, we will restrict our attention to spacetimes with homogeneous (time-dependent) conformal factors. A homogeneous conformal transformation is defined as a conformal transformation with a conformal factor that satisfies
$\gamma_{\mu}^{\alpha}\nabla_{\alpha}\Omega=0.$ (14)
This means that the conformal factor does not vary in spatial directions, and therefore only depends on time when expressed in suitable coordinates. As will be seen later, the time-dependent scale factor used to represent expanding cosmologies is one example of a homogeneous conformal factor.
First, we need to determine the conformal weights of the quantities contributing to the polarization wiggle rate. To do so, we need the relationship between the Christoffel symbols of the two geometries [24]:
${\tilde{\Gamma}}_{\mu\nu}^{\rho}=\Gamma_{\mu\nu}^{\rho}+C_{\mu\nu}^{\rho},$ (15)
where
$C_{\mu\nu}^{\rho}\equiv\Omega^{-1}(\delta_{\mu}^{\rho}\nabla_{\nu}\Omega+\delta_{\nu}^{\rho}\nabla_{\mu}\Omega-g_{\mu\nu}g^{\rho\lambda}\nabla_{\lambda}\Omega).$ (16)
${\tilde{\Gamma}}_{\mu\nu}^{\rho}$ and  $\Gamma_{\mu\nu}^{\rho}$ denote the Christoffel symbols of the physical and conformal frame geometries, respectively.
III.1 Conformal Properties of Relevant Quantities
It is known that, in 4 dimensions, Maxwell’s theory of electromagnetism in vacuum is conformally invariant. A consequence of this invariance is that the vector potential and 4-momentum transform as follows under a conformal transformation with conformal factor $\Omega^{2}$:
${\tilde{A}}_{\mu}=A_{\mu},{\tilde{A}}^{\mu}=\Omega^{-2}A^{\mu}$ (17)
${\tilde{p}}_{\mu}=p_{\mu},{\tilde{p}}^{\mu}=\Omega^{-2}p^{\mu}.$ (18)
Then, let us consider the rest frame geodesic congruence in the physical frame, ${\tilde{u}}^{\mu}$. Define a congruence $u^{\mu}=\Omega{\tilde{u}}^{\mu}$ in the conformal frame. $u^{\mu}$ is evidently a timelike unit vector in the conformal frame, but it is not apparent that it also satisfies the geodesic equation. Since ${\tilde{u}}^{\mu}$ by definition is a geodesic in the physical frame, it satisfies the geodesic equation ~${\tilde{\nabla}}_{\tilde{u}}{\tilde{u}}^{\mu}=0$. By the use of eqs. (15) and (16), this equation may be expanded in terms of $u^{\mu}$, giving the equation
$u^{\mu}\nabla_{\mu}u_{\nu}+\Omega^{-1}\gamma_{\nu}^{\mu}\nabla_{\mu}\Omega=0.$
Since the conformal factor is homogeneous, we may use eq. (14). We then get
$u^{\mu}\nabla_{\mu}u_{\nu}=0.$
Thus, $u^{\mu}=\Omega{\tilde{u}}^{\mu}$ is a rest frame geodesic in the conformal frame $\mathcal{M}$. Notice that this conformal relation is not valid for an arbitrary timelike geodesic in $\tilde{\mathcal{M}}$. It is only valid for the rest frame geodesic ${\tilde{u}}^{\mu}$.
Next, given any null geodesic ${\tilde{k}}^{\mu}$ in the physical frame, we find that $k^{\mu}=\Omega^{2}{\tilde{k}}^{\mu}$ satisfies
$\nabla_{k}k^{\mu}=0.$
This implies that $k^{\mu}$ is a null geodesic in the conformal frame. The conformal weights of the quantities that depend on the two geodesics ${\tilde{p}}^{\mu}$ and ${\tilde{u}}^{\mu}$ can now be determined:
${\tilde{u}}_{\mu}=\Omega u_{\mu},{\tilde{u}}^{\mu}=\Omega^{-1}u^{\mu}$ (19)
$\tilde{p}=\Omega^{-1}p$ (20)
${\tilde{\hat{p}}}_{\mu}=\Omega{\hat{p}}_{\mu},{\tilde{\hat{p}}}^{\mu}=\Omega^{-1}{\hat{p}}^{\mu}$ (21)
${\tilde{S}}_{\mu\nu}=\Omega^{2}S_{\mu\nu},{\tilde{S}}^{\mu\nu}=\Omega^{-2}S^{\mu\nu}$ (22)
${\tilde{\epsilon}}_{\mu\nu}=\Omega^{2}\epsilon_{\mu\nu},{\tilde{\epsilon}}^{\mu\nu}=\Omega^{-2}\epsilon^{\mu\nu}.$ (23)
The transverse polarization vector $\epsilon^{\mu}$ transforms as a unit vector:
${\tilde{\epsilon}}_{\mu}=\Omega\epsilon_{\mu},{\tilde{\epsilon}}^{\mu}=\Omega^{-1}\epsilon^{\mu}.$
The polarization degree $\mathcal{P}$ and the circular polarization degree $\mathcal{V}$ are invariant under conformal transformations:
$\tilde{\mathcal{P}}=\mathcal{P}$ (24)
$\tilde{\mathcal{V}}=\mathcal{V}.$ (25)
III.2 Conformal Transformation of the Curvature Twist
The Weyl tensor transforms as follows under conformal transformations [24]:
${{\tilde{C}}^{\rho}}_{\sigma\mu\nu}={C^{\rho}}_{\sigma\mu\nu}.$ (26)
Using eqs. (19)-(26), we obtain the conformal transformation rule of the curvature twist, as expressed by eq. (7):
$\tilde{\mathcal{Z}}=\Omega^{-2}Z.$ (27)
Eq. (27) is remarkably simple. It complies with the intuitive notion that, if you scale a geometry by a fixed, coordinate-independent scale factor $\Omega$, the components of the Riemann tensor would be scaled by a factor $\Omega^{-2}$.
III.3 The Scale Dependency of the Polarization Wiggle Rates
It is now straight forward to determine how the polarization wiggle rates $\omega$ and $\chi$ behave under homogeneous conformal transformations. From eqs. (18) and (27) follows that, if $\omega/p$ and $\chi/p$ are solutions to eqs. (8) and (9) in the conformal frame, so are $\tilde{\omega}/\tilde{p}$ and $\tilde{\chi}/\tilde{p}$. This implies that, given the same initial conditions,
$\frac{\tilde{\omega}}{\tilde{p}}=\frac{\omega}{p}$
$\frac{\tilde{\chi}}{\tilde{p}}=\frac{\chi}{p}.$
By using eq. (20), we find the polarization wiggle rates in the physical frame in terms of the conformal frame values:
$\tilde{\omega}=\Omega^{-1}\omega$ (28)
$\tilde{\chi}=\Omega^{-1}\chi.$ (29)
Eqs. (28) and (29) are generally valid for any spacetime with an arbitrary 4-dimensional conformal frame geometry and a homogeneous (time-dependent) conformal factor.
IV Polarization Wiggling in Linear Gravity
In the following, we will apply linear gravity to analyze the polarization wiggling effect in spacetimes with conformally flat background geometries. The physical metric $\tilde{g}$ can be written
${\tilde{g}}_{\mu\nu}={a(\eta)}^{2}g_{\mu\nu}={a(\eta)}^{2}\left(\eta_{\mu\nu}+h_{\mu\nu}\right),$ (30)
where $g_{\mu\nu}\equiv\eta_{\mu\nu}+h_{\mu\nu}$ denotes the conformal frame metric, $\eta_{\mu\nu}$ is the Minkowski metric, $h_{\mu\nu}$ a metric perturbation in the conformal frame, $\eta$ denotes conformal time (time in the conformal frame) and $a(\eta)$ is a homogeneous (time dependent) scale factor.
By applying the results of the previous section, we find that this way of relating the physical spacetime to a geometry with a flat background through a conformal transformation allows us to study polarization wiggling in the presumably much simpler conformal frame geometry and use the results of the previous section to transform the results to the physical frame. Given that the conformal frame geometry $g$ is derived from a physical frame geometry $\tilde{g}$ that satisfies the Einstein equations, this approach offers significant simplifications to the analysis of polarization wiggling in the physical frame. For instance, by applying eq. (28), the polarization axis wiggle rate $\tilde{\omega}$ in the physical frame can be determined from the quantity in the conformal frame:
$\tilde{\omega}(\eta)={a(\eta)}^{-1}\omega(\eta).$ (31)
In linear gravity with a Minkowski background, the linearized Riemann tensor is [24]
$R_{\mu\sigma\alpha\lambda}=\frac{1}{2}\left(h_{\mu\lambda,\sigma\alpha}+h_{\sigma\alpha,\mu\lambda}-h_{\mu\alpha,\sigma\lambda}-h_{\sigma\lambda,\mu\alpha}\right)+\mathcal{O}(h^{2}).$ (32)
The curvature twist derived from eq. (32) is
$\mathcal{Z}=u^{\left[\mu\right.}{\hat{p}}^{\left.\alpha\right]}\epsilon^{\beta\lambda}h_{\alpha\beta,\lambda\mu}.$ (33)
By the use of eq. (3), eq. (33) can be rewritten as follows:
$\mathcal{Z}=\frac{1}{p}\left(\nabla_{u}\nu^{\left(p\right)}-\nabla_{p}\nu^{\left(u\right)}\right)=\nabla_{u}\nu^{\left(\hat{p}\right)}-\nabla_{\hat{p}}\nu^{\left(u\right)},$ (34)
where
$\nu^{\left(q\right)}\equiv\frac{1}{2}\epsilon^{\sigma\lambda}q^{\alpha}h_{\alpha\sigma,\lambda}$ (35)
for an arbitrary 4-vector $q^{\mu}$. While the null geodesic is parametrized by the affine parameter $\lambda$, the timelike rest-frame geodesics are parametrized by proper time $\tau$. Then, $\nabla_{p}\nu^{(u)}=d\nu^{(u)}/d\lambda$, while $\nabla_{u}\nu^{(p)}=d\nu^{(p)}/d\tau$. Using that to lowest order, $d\nu^{(p)}/d\tau=\partial\nu^{(p)}/\partial\eta$, eq. (34) can be approximated as
$\mathcal{Z}=\frac{1}{p}\left(\frac{\partial}{\partial\eta}\nu^{\left(p\right)}-\frac{d\nu^{\left(u\right)}}{d\lambda}\right)+\mathcal{O}(h^{2}).$ (36)
The second term on the right-hand sidef of eq. (36) can readily be integrated, and using this expression for the curvature twist in eqs. (12) and (13) gives the following expressions for the wiggle rates:
$\omega(\lambda)=\omega(\lambda_{*})+\operatorname*{\int}\limits_{\lambda_{*}}^{\lambda}d\lambda^{\prime}\frac{\partial\nu^{\left(p\right)}}{\partial\eta}-{\Delta\nu}^{\left(u\right)}$ (37)
$\chi(\lambda)=-\mathcal{V}\omega(\lambda)+\chi(\lambda_{*})+\mathcal{V}\omega(\lambda_{*}),$ (38)
where
${\Delta\nu}^{\left(u\right)}\equiv\nu^{\left(u\right)}(\lambda)-\nu^{\left(u\right)}(\lambda_{*}).$
In eqs. (37) and (38), we used that to lowest order with a Minkowski background, $p(\lambda)/p(\lambda_{*})=1+\mathcal{O}(h)$.
V Scalar-Vector-Tensor Decomposition
In linear gravity, the metric perturbation $h_{\mu\nu}$ can be decomposed into irreducible scalar, vector and tensor components [25, 26, 27]. This means that, in any metric gauge, we may separate an arbitrary metric perturbation $h_{\mu\nu}$ into independent scalar, vector and tensor perturbations:
$h_{\mu\nu}=h_{\mu\nu}^{\left(S\right)}+h_{\mu\nu}^{\left(V\right)}+h_{\mu\nu}^{\left(T\right)}.$ (39)
Since scalar, vector and tensor perturbations are sourced and evolve independently in linear gravity, their effects can be analyzed separately. Therefore, since the curvature twist, as expressed by eq. (33), is linear in the metric perturbation $h_{\mu\nu}$, we may split both the curvature twist and polarization wiggle rates into scalar, vector and tensor terms that are induced by corresponding scalar, vector and tensor perturbations to the metric:
$\mathcal{Z}=Z^{\left(S\right)}+Z^{\left(V\right)}+Z^{\left(T\right)}$
$\omega=\omega^{\left(S\right)}+\omega^{\left(V\right)}+\omega^{\left(T\right)},$
where
$\mathcal{Z}^{\left(Q\right)}\equiv u^{\left[\mu\right.}{\hat{p}}^{\left.\alpha\right]}\epsilon^{\beta\lambda}h_{\alpha\beta,\lambda\mu}^{\left(Q\right)}$ (40)
$\omega^{\left(Q\right)}(\lambda)\equiv p\text{}\operatorname*{\int}\limits_{\lambda_{*}}^{\lambda}d\lambda^{\prime}\mathcal{Z}^{\left(Q\right)}(\lambda^{\prime})$ (41)
for tensorial order $Q=S,V,T$.
V.1 Gauge Invariance of Curvature Twist
The curvature twist $\mathcal{Z}$ is by definition a scalar quantity, which means it is gauge invariant in linear gravity for an arbitrary metric perturbation $h_{\mu\nu}$. Since the metric perturbations $h_{\mu\nu}^{(S)},h_{\mu\nu}^{(V)}$ and $h_{\mu\nu}^{(T)}$ can be treated as independent perturbations, the corresponding curvature twist components $\mathcal{Z}^{(S)},\mathcal{Z}^{(V)}$ and $\mathcal{Z}^{(T)}$ are individially gauge invariant. This implies that the corresponding polarization wiggle rates $\omega^{(S)},\omega^{(V)}$ and $\omega^{(T)}$ are also gauge invariant. This is an important finding, because it means that we have found a way to physically quantify the effects of scalar, vector and tensor perturbations independently. Furthermore, we are free to choose metric gauge to evaluate the curvature twist and polarization wiggle rate components.
Next, we will choose a particular metric gauge and evaluate $\mathcal{Z}^{(S)},\mathcal{Z}^{(V)}$ and $\mathcal{Z}^{(T)}$ individually in the chosen gauge.
V.2 Metric Decomposition
Let us do a scalar-vector-tensor decomposition of the metric perturbation $h_{\mu\nu}$into irreducible scalar vector and tensor components. We will apply transverse gauge. In this gauge the metric takes the form [24]
$h_{00}=h_{00}^{\left(S\right)}=-2\psi$
$h_{0i}=h_{0i}^{\left(V\right)}=w_{i}$
$h_{\mathrm{ij}}=h_{\mathrm{ij}}^{\left(S\right)}+h_{\mathrm{ij}}^{\left(T\right)}=2\phi\delta_{\mathrm{ij}}+h_{\mathrm{ij}}^{\mathrm{TT}},$
where $\psi$ and $\phi$ are two scalar potentials, $w^{i}$ is a transverse vector and $h_{\mathrm{ij}}^{\mathrm{TT}}$ is a transverse and traceless tensor.
VI Scalar-induced Polarization Wiggling
In the transverse gauge, the scalar component of the curvature twist is
$\mathcal{Z}^{\left(S\right)}={\hat{p}}^{j}{\hat{p}}^{l}\varepsilon^{\mathrm{ljk}}{\phi}_{,\mathrm{k0}}.$
Since ${\hat{p}}^{j}{\hat{p}}^{l}\varepsilon^{\mathrm{ljk}}=0$, we find that $\mathcal{Z}^{(S)}=0$. Consequently,
$\omega^{\left(S\right)}=0.$ (42)
Since $\mathcal{Z}^{(S)}$ is gauge invariant, we can conclude that scalar perturbations to the metric do not contribute to polarization wiggling.
VII Vector-induced Polarization Wiggling
VII.1 Polarization Wiggling from the Gravitomagnetic Field
In the transverse gauge, the vector component of the curvature twist is
$\mathcal{Z}^{\left(V\right)}=\frac{1}{2}\hat{\text{\boldmath{\hbox{p}}}}\cdot\text{\boldmath{\hbox{\nabla}}}\left(\hat{\text{\boldmath{\hbox{p}}}}\cdot\text{\boldmath{\hbox{\nabla}}}\times\text{\boldmath{\hbox{w}}}\right)$
[…]
$\mathcal{Z}^{\left(V\right)}=\frac{1}{2}\hat{\text{\boldmath{\hbox{p}}}}\cdot\text{\boldmath{\hbox{\nabla}}}\mathcal{B}$ (43)
where $\mathcal{B}\equiv\hat{\text{\boldmath$p$}}\cdot\text{\boldmath$\mathcal{B}$}$. We will refer to $\mathcal{B}$ as the longitudinal gravitomagnetic field. The projection of the gravitomagnetic field $\mathcal{B}$ along the direction of propagation $\hat{p}$ induces a  rotation of the polarization axis of the radiation as it propagates along its null geodesic. This effect of the gravitomagnetic field on polarized electromagnetic radiation is referred to as the Gravitational Faraday Effect, due to its analogy with the Faraday Effect, which is the effect an ambient magnetic field has on the polarization of passing electromagnetic radiation. In other words, the Gravitational Faraday Effect can be viewed as polarization wiggling induced by vector perturbations through the resulting gravitomagnetic field.
The vector-induced polarization wiggle rate, as given by eq. (41), can then be expressed in terms of the spatial gradient of the longitudinal gravitomagnetic field:
$\omega^{\left(V\right)}(\lambda)\equiv\frac{1}{2}p\text{}\operatorname*{\int}\limits_{\lambda_{*}}^{\lambda}d\lambda^{\prime}\hat{\text{\boldmath{\hbox{p}}}}\cdot\text{\boldmath{\hbox{\nabla}}}\mathcal{B}.$
Thus, polarization wiggling can only be induced if the longitudinal gravitomagnetic field possesses a non-zero spatial gradient along the direction of propagation of the polarized radiation.
Alternatively, we may use eq. (36). The terms $\nu^{(u)}$ and $\nu^{(p)}$ evaluate to
$\nu^{\left(p\right)}=-\frac{p}{2}\hat{\text{\boldmath$p$}}\cdot\text{\boldmath$\nabla$}\times\text{\boldmath$w$}=-\frac{1}{2}p\mathcal{B}$ (44)
$\nu^{\left(u\right)}=-\frac{1}{2}\hat{\text{\boldmath$p$}}\cdot\text{\boldmath$\nabla$}\times\text{\boldmath$w$}=-\frac{1}{2}\mathcal{B}$ (45)
The vector component of the curvature twist then evaluates to
$\mathcal{Z}^{\left(V\right)}=-\frac{1}{2}\left(\frac{\partial\mathcal{B}}{\partial\eta}-\frac{1}{p}\frac{d\mathcal{B}}{d\lambda}\right),$
where we used that  $\frac{d\mathcal{B}}{d\tau}=\frac{\partial\mathcal{B}}{\partial\eta}$ to first order in the metric perturbation. By eq. (37), the polarization axis wiggle rate then is
$\omega^{\left(V\right)}(\lambda)=-\frac{1}{2}\left(p\operatorname*{\int}\limits_{\lambda_{*}}^{\lambda}d\lambda^{\prime}\frac{\partial\mathcal{B}}{\partial\eta}-\Delta\mathcal{B}\right),$ (46)
where
$\Delta\mathcal{B}\equiv\mathcal{B}(\lambda)-\mathcal{B}(\lambda_{*}).$
We notice that, even if the gravitomagnetic field is stationary with $\partial\mathcal{B}/\partial\eta=0$, the vectorial contribution to the polarization axis wiggle rate, $\omega^{(V)}$, is still non-zero if the change in the longitudinal gravitomagnetic field between the emitter and detector is non-zero. The term $\Delta\mathcal{B}$ on the right-hand side of eq. (46) stems from the term $\nu^{(u)}$, and can be ascribed to the difference in inertial frame-dragging rate between the emitter and detector frames. This can be seen as follows: The angular velocity of the local inertial frame as measured by a distant observer in a rest frame is [28]
${\text{\boldmath{\hbox{\omega}}}}_{f}=-\frac{1}{2}\text{\boldmath{\hbox{\nabla}}}\times\text{\boldmath{\hbox{w}}}=-\frac{1}{2}\text{\boldmath{\hbox{\mathcal{B}}}}$
This  rotation of the local inertial frame is the manifestation of a gravitational effect known as the frame dragging effect, or dragging of inertial frames. For an emitter of radiation with polarization that is constant in the frame of emission, the polarization axis will rotate with the frame. The natural interpretation of eq. (46) is therefore that $\omega^{(V)}$ is the net angular speed of the polarization axis in the frame of observation after subtracting the difference $\Delta\mathcal{B}$ in angular frame dragging rates between the inertial frames of observation and emission.
Let us now consider eq. (46) for a few special cases. First of all, an interesting consequence of the result of eq. (46) is that, given a distant observer at rest observing an inertial emitter of polarized electromagnetic radiation with constant polarization orbiting a gravitational source with a stationary gravitomagnetic field, the polarization axis wiggle rate of the observed polarized radiation is a direct measure of the angular frame dragging rate of the frame of emission:
$\omega^{\left(V\right)}(\lambda)=-\frac{1}{2}\mathcal{B}(\lambda_{*})=\omega_{f}(\lambda_{*}),$ (47)
where $\omega_{f}\equiv\hat{\text{\boldmath$p$}}\cdot{\text{\boldmath$\omega$}}_{f}$ is the frame dragging rate around the direction of propagation. Conversely, given a distant emitter of polarized electromagnetic radiation with constant polarization at a position with no ambient gravitomagnetic field and a detector moving inertially in the vicinity of a gravitational source with a stationary gravitomagnetic field, the polarization axis wiggle rate of the measured polarized radiation is a direct measure of the angular frame dragging rate of the frame of observation:
$\omega^{\left(V\right)}(\lambda)=\frac{1}{2}\mathcal{B}(\lambda)=-\omega_{f}(\lambda).$ (48)
Finally, consider  eq. (46) for the case $\Delta\mathcal{B}=0$, which occurs when the gravitomagnetic field is negligible at both the detector and emitter locations. In this case, $\omega^{(V)}$ is non-zero only if the radiation traverses the vicinity of a gravitational source with a non-stationary gravitomagnetic field. The gravitomagnetic field is sourced by the matter current $j^{i}\equiv T^{0i}\simeq\rho v^{i}$ [29], where $T^{\mu\nu}$ is the energy-momentum tensor, $\rho$ is matter density and $v^{i}$ is matter velocity. Therefore, changing matter currents induce changes to the gravitomagnetic field, which in turn induce polarization wiggling in any passing electromagnetic radiation.
We conclude that the polarization axis wiggle rate $\omega^{(V)}$ is a direct measure of the longitudinal gravitomagnetic field $\mathcal{B}$. The contribution to the polarization axis wiggle rate from vector perturbations has two terms. The first contribution is due to the difference in the angular frame dragging rate between the frame of emission and the frame of observation. The second contribution comes from the presence of non-stationary gravitomagnetic fields along the radiation trajectory.
It is worth remarking that the polarization wiggle rate induced by a localized gravitomagnetic field on a Minkowski backgrond is to a large extent insensitive to distance between the emitter and detector. This can be seen from eq. (46). The second term on the right-hand side is clearly independent of distance, while the first term is independent of distance because $\mathcal{B}$ is non-zero in just a limited part of the path between emitter and detector. This characteristic distinguishes the polarization wiggle rate as a radiation observable from radiation intensity, which typically decreases as the radiation spreads with distance.
VII.2 Measuring Angular Momentum of a Gravitational Source using Polarization Wiggling
We will now apply the results for arbitrary scalar and vector perturbations to analyze a more specific case. We will show that the polarization wiggling of electromagentic radiation emitted by a radiation emitter in orbit around a gravitational source with unknown angular momentum can be used to determine the angular momentum of the gravitational source.
Let us begin by stating the spacetime metric in the vicinity of the gravitational source. We will assume it to be a stationary axi-symmetric metric with a line element [30]
${\mathrm{ds}}^{2}=-\left(1-\frac{2m}{r}\right){\mathrm{dt}}^{2}+\left(1+\frac{2m}{r}\right)\left({\mathrm{dr}}^{2}+r^{2}{\mathrm{d\Omega}}^{2}\right)-\frac{4j{\sin}^{2}\theta}{r}\mathrm{dt}\mathrm{d\phi},$ (49)
where $t$ is time, $(r,\theta,\phi)$ are spherical coordinates, $m$ is the mass of the gravitational source and $j$ is the scalar angular momentum, assuming the rotation axis is the z-axis $\hat{z}$. In Section VI, we found that scalar perturbations do not contribute to polarization wiggling. The only component of the above metric that contributes is the $\mathrm{dt}\mathrm{d\phi}$ term. The vector perturbation is therefore
$w_{r}=w_{\theta}=0,w_{\phi}=-\frac{4j{\sin}^{2}\theta}{r}.$
The gravitomagentic field $\mathcal{B}$ is then
B×$\text{\boldmath{\hbox{\mathcal{B}}}}\equiv\text{\boldmath{\hbox{\nabla}}}\times\text{\boldmath{\hbox{w}}}=-\frac{12j}{r^{2}}\mu\sqrt{1-\mu^{2}}\text{\boldmath{\hbox{n}}}\text{\boldmath{\hbox{,}}}$ (50)
[…]
VIII Tensor-induced Polarization Wiggling
The tensor perturbation $h_{\mu\nu}^{(T)}$ of eq. (39) is a gauge invariant quantity that is transverse and traceless, conventionally denoted $h_{\mu\nu}^{\mathrm{TT}}$. To give the results of the previous sections a proper demonstration, we will now explore the the polarization wiggle rate $\tilde{\omega}$ arising from a single gravitational tensor mode in linear gravity, assuming a spacetime with a conformally flat background. This will cover spacetimes with both a flat Minkowski background as well as expanding cosmologies.
Let us assume a cosmology with a conformally flat background and the metric of eq. (30) with a transverse, traceless metric perturbation $h_{\mu\nu}^{\mathrm{TT}}$. The spatial Einstein equation in this case is [31]
${\left(h_{\mu\nu}^{\mathrm{TT}}\right)}^{\prime\prime}+2{\mathcal{H}(h_{\mu\nu}^{\mathrm{TT}})}^{\prime}-\nabla^{2}h_{\mu\nu}^{\mathrm{TT}}=0,$ (52)
where $\mathcal{H}\equiv a^{\prime}/a$ is the conformal Hubble factor and ‘ denotes differentiation with respect to conformal time $\eta$. We will not attempt to study the effect of a general tensor perturbation $h_{\mu\nu}^{\mathrm{TT}}$. Instead, under the presumption that an arbitrary tensor perturbation $h_{\mu\nu}^{\mathrm{TT}}$ can be represented as a linear combination of simple plane wave components, we will study the effect of a single plane gravitational wave (tensor mode) with arbitrary polarization. The polarization wiggle rate indued by a composite gravitational tensor perturbation can then simply be aggregated as the sum of the polarization wiggle rates of the plane wave components of the tensor perturbation.
VIII.1 Wiggle Rate from a Tensor Mode with a Conformally Flat Background
We will now explore the polarization wiggle rate $\tilde{\omega}$ in the physical frame arising from a single gravitational tensor mode propagating in an expanding universe with a conformally flat background geometry. Let us apply the following trial solution to eq. (52):
$h_{\mu\nu}^{\mathrm{TT}}=C_{\mu\nu}(\eta)e^{ik{\hat{k}}_{\mu}x^{\mu}},$ (53)
where $\hat{k}$ is a spatial unit vector, and $k>0$ is an arbitrary angular frequency. Notice that we use a complex representation of the tensor mode for the sake of analytic and notational simplicity, with the implicit understanding that only the real part of expressions derived linearly from it should be used in physical quantities. ${\hat{k}}^{\mu}$ is the direction of propagation, and $k$ is the angular frequency of the gravitational mode. $C_{\mu\nu}(\eta)$ is a time-dependent polarization tensor.
Eq. (53) is a solution to eq. (52)  if $C_{\mu\nu}$ satisfies
$C_{\mu\nu}^{\prime\prime}+2\mathcal{H}C_{\mu\nu}^{\prime}+k^{2}C_{\mu\nu}=0.$ (54)
[…]
$C_{\mu\nu}(\eta)=\sum\limits_{s=+,\times}\sigma_{\mu\nu}^{s}(\hat{k})c_{s}(k,\eta),$
where the sum is over polarizations $s=+,\times$. $c_{s}(k,\eta)$ is the amplitude of the polarization $s$ component of the tensor mode. The basis tensors $\sigma_{\mu\nu}^{s}(\hat{k})$ are assumed traceless, and they must satisfy $u^{\mu}\sigma_{\mu\nu}^{s}={\hat{k}}^{\mu}\sigma_{\mu\nu}^{s}=0$ and the normalization condition $\sigma_{\mu\nu}^{s}(\hat{k})\sigma^{s^{\prime}\mu\nu}(\hat{k})=2\delta^{ss^{\prime}}$ [31]. This ensures that the metric perturbation of eq. (53) is transverse and traceless by construction.
From eq. (54) follows that the tensor mode amplitudes must satisfy the differential equation
$c_{s}^{\prime\prime}+2\mathcal{H}c_{s}^{\prime}+k^{2}c_{s}=0.$ (55)
First, let us try to determine the amplitude $c_{s}(\eta)$ by solving eq. (55) with the following trial solution:
$c_{s}(k,\eta)=h_{s}(k,\eta)e^{-ik\eta}.$ (56)
Eq. (55) then takes the form
$e^{ik\eta}{a^{-2}(a^{2}(h_{s}^{\prime}-ikh_{s}))}^{\prime}=0,$
which has the solution
$a^{2}(h_{s}^{\prime}-ikh_{s})=-ikh_{s}^{0}=\mathrm{const},$ (57)
where $h_{s}^{0}$ is a constant of integration. Let us assume that $h_{s}(\eta)$ is varying slowly, with $|h_{s}^{\prime}|\ll|kh_{s}|$. From eq. (57) then follows that
$h_{s}(k,\eta)\simeq\frac{h_{s}^{0}(k)}{{a(\eta)}^{2}}.$ (58)
We have now established the time dependency of the polarization amplitude $c_{s}(k,\eta)$ for a general tensor mode. Next, let us evaluate the polarization wiggle rate $\omega$ in the conformal frame.
From eq. (41) follows that the polarization wiggle rate is
$\omega^{\left(T\right)}(\lambda)\equiv p\text{}\operatorname*{\int}\limits_{\lambda_{*}}^{\lambda}d\lambda^{\prime}\mathcal{Z}^{\left(T\right)}(\lambda^{\prime}).$ (59)
For background null geodesics in the conformal frame, we have $d\eta/d\lambda=p=\mathrm{const}$. We may use this to reparametrize  both the radiation and the gravitational mode geodesics in terms of conformal time. Eq. (59) can then be written
$\omega^{\left(T\right)}(\eta)=\operatorname*{\int}\limits_{\eta_{*}}^{\eta}d\eta^{\prime}\mathcal{Z}^{\left(T\right)}(\eta^{\prime}),$ (60)
where $\eta_{*}$ is the time of emission.
We will now evaluate the wiggle rate $\omega$ in the conformal frame by using eq. (60). The wiggle rate $\tilde{\omega}$ in the physical frame can then be derived by using eq. (31). Since we are evaluating $\mathcal{Z}^{(T)}$ to first order in the metric perturbation, we may use the geodesic in the Minkowski background when evaluating it. Assume that the observer is at the origin with a conformal frame rest frame geodesic $x_{0}^{\mu}(\eta)=\eta u^{\mu}$. Let the radiation geodesic in the conformal frame be $x^{\mu}(\lambda)=x_{0}^{\mu}(\eta)+p^{\mu}(\lambda-\lambda_{0})$, where $\lambda_{0}$ labels the observation event. Then
${\hat{k}}_{\mu}x^{\mu}=-p\mu(\lambda_{0}-\lambda)=-\mu(\eta-\eta^{\prime}),$
where $\mu\equiv{\hat{k}}_{\mu}{\hat{p}}^{\mu}$, $\eta$ is the time of observation and $\eta^{\prime}$ is the time parameter of an arbitrary point along the radiation geodesic. $\mu=\cos\theta$, where $\theta$ is the angle of inclination of the observed radiation with the propagation direction of the tensor mode.
Using eq. (40), the curvature twist in the conformal frame of this metric is
$\mathcal{Z}^{\left(T\right)}(\eta^{\prime})=ik{\hat{p}}^{\mu}\epsilon^{\nu\lambda}{\hat{k}}_{\lambda}C_{\mu\nu}^{\prime}e^{ik{\hat{k}}_{\beta}x^{\beta}}=ik\sum\limits_{s}\sigma^{s}(\hat{k},\hat{p})c_{s}^{\prime}e^{-ik\mu(\eta-\eta^{\prime})},$ (61)
where we have introduced the angular function
$\sigma^{s}(\hat{k},\hat{p})\equiv{\hat{p}}^{\mu}\sigma_{\mu\nu}^{s}(\hat{k})l^{\nu},$ (62)
with $l^{\nu}\equiv\epsilon^{\nu\lambda}{\hat{k}}_{\lambda}$. We notice that the vector $l^{\nu}$ is transverse to both $\hat{p}$ and $\hat{k}$. The scalar $\sigma^{s}(\hat{k},\hat{p})$ represents the coupling between the polarization $s$ component of a gravitational tensor mode propagating in direction $\hat{k}$ and the polarization of electromagnetic radiation propagating in direction $\hat{p}$. We will refer to $\sigma^{s}(\hat{k},\hat{p})$ as the gravitational polarization sensitivity of polarized radiation emitted in direction $\hat{p}$ to a gravitational tensor mode with polarization $s$ propagating in direction $\hat{k}$. It is a purely geometric quanty that will be analyzed in detail in the next section. The right-hand side of eq. (60) can be integrated by parts, and we obtain
$\omega^{\left(T\right)}(\eta)=\mathrm{Re}\left\{ik\sum\limits_{s}\sigma^{s}(\hat{k},\hat{p})\left(c_{s}(k,\eta)-c_{s}(k,\eta_{*})e^{-ik\mu(\eta-\eta_{*})}-ik\mu e^{-ik\mu\eta}\operatorname*{\int}\limits_{\eta_{*}}^{\eta}d\eta^{\prime}e^{ik\mu\eta^{\prime}}c_{s}(k,\eta^{\prime})\right)\right\}.$ (63)
Inserting the $c_{s}$ solution of eqs. (56) and (58) into eq. (63) yields the following expression for the polarization wiggle rate:
$\omega^{\left(T\right)}(\eta)=\mathrm{Re}\left\{ik\left(e^{-ik\eta}-e^{-ik\mu\eta}\left(\frac{e^{-ik(1-\mu)\eta_{*}}}{{a(\eta_{*})}^{2}}+ik\mu\operatorname*{\int}\limits_{\eta_{*}}^{\eta}d\eta^{\prime}\frac{e^{-ik\left(1-\mu\right)\eta^{\prime}}}{{a(\eta^{\prime})}^{2}}\right)\right)\sum\limits_{s}\sigma^{s}(\hat{k},\hat{p})h_{s}^{0}(k)\right\}.$ (64)
Assuming that both the detector and emitter follow rest geodesics, we may assume that the comoving distance between them is constant to lowest order:
$\chi_{*}\equiv\eta-\eta_{*}=\mathrm{const}.$
The wiggle rate can then be expressed in terms of comoving distance:
$\omega^{\left(T\right)}(\eta)=\mathrm{Re}\left\{ike^{-ik\eta}\left(1-\frac{e^{i\left(1-\mu\right)k\chi_{*}}}{{a(\eta-\chi_{*})}^{2}}-ik\mu\ \ \operatorname*{\int}\limits_{0}^{\chi_{*}}d\chi\frac{e^{ik\left(1-\mu\right)\chi}}{{a(\eta-\chi)}^{2}}\right)\sum\limits_{s}\sigma^{s}(\hat{k},\hat{p})h_{s}^{0}(k)\right\}.$ (65)
Eq. (64) gives the polarization wiggle rate $\omega$ in the conformal frame for any cosmology with a conformally flat background. Furthermore, since at present time $\eta=\eta_{0}$, $a(\eta_{0})=1$, the present time polarization wiggle rate in the physical frame, which is given by eq. (31), is equal to the conformal frame value:
${\tilde{\omega}}^{\left(T\right)}(\eta_{0})=\omega^{\left(T\right)}(\eta_{0}).$
It is apparent from eq. (64) that, if $a(\eta_{*})\ll 1$, the second term on the right-hand side will be the dominant one of the two first terms. This is the case if the polarized radiation and the tensor mode were emitted at early times in an expanding cosmology. In the next two sections, we will evaluate eq. (64) for two different backgrounds: A flat Minkowski background and an expanding, conformally flat background.
VIII.2 Gravitational Polarization Sensitivity $\sigma^{s}$
Next, let us evaluate the polarization sensitivity $\sigma^{s}(\hat{k},\hat{p})$ of eq. (62). It is defined as a purely geometric quantity in terms of two spatial directions $\hat{k}$ and $\hat{p}$ and the tensor polarization basis $\sigma_{\mu\nu}^{s}$. First, the tensor basis $\sigma_{\mu\nu}^{s}(\hat{k})$ for a given direction of propagation ${\hat{k}}^{\mu}$ must be chosen. With the choice of an orthonormal polarization basis $e_{A}^{\mu}(\hat{k}),A=1,2$ perpendicular to the propagation direction $\hat{k}$, the basis tensor $\sigma_{\mu\nu}^{s}$ can be defined by
$\sigma_{\mu\nu}^{s}(\hat{k})\equiv\sigma_{\mathrm{AB}}^{s}e_{\mu}^{A}(\hat{k})e_{\nu}^{B}(\hat{k}),$
where $\sigma_{\mathrm{AB}}^{s}$ are the two symmetric Pauli matrices
$\sigma_{\mathrm{AB}}^{+}\equiv\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right),\sigma_{\mathrm{AB}}^{\times}\equiv\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right).$
Then,
$\sigma_{\mu\nu}^{+}=e_{\mu}^{1}e_{\nu}^{1}-e_{\mu}^{2}e_{\nu}^{2}$ (66)
$\sigma_{\mu\nu}^{\times}=e_{\mu}^{1}e_{\nu}^{2}+e_{\mu}^{2}e_{\nu}^{1}.$ (67)
Using the definition of the screen rotator in eq. (5),  $\sigma^{s}$, as defined in eq. (62), can be expressed as
$\sigma^{+}(\hat{k},\hat{p})=-{\hat{p}}^{\mu}\sigma_{\mu\nu}^{\times}{\hat{p}}^{\nu}$
$\sigma^{\times}(\hat{k},\hat{p})={\hat{p}}^{\mu}\sigma_{\mu\nu}^{+}{\hat{p}}^{\nu}.$
Furthermore, using eqs. (66) and (67), we get
$\sigma^{+}(\hat{k},\hat{p})=-2\beta_{1}\beta_{2}$ (68)
$\sigma^{\times}(\hat{k},\hat{p})=\beta_{2}^{2}-\beta_{1}^{2}$ (69)
where $\beta_{A}(\hat{k},\hat{p})\equiv{\hat{p}}_{\mu}e_{A}^{\mu}(\hat{k}),A=1,2$ are the projections of the radiation propagation direction $\hat{p}$ against the polarization basis vector $e_{A}(\hat{k})$ for direction $\hat{k}$.
VIII.3 Wiggle Rate from a Tensor Mode on a Flat Background
Let us assume a flat background with $a(\eta)=\mathrm{const}=1$ and evaluate the wiggle rate of eq. (64). The result is
$\omega^{\left(T\right)}(\eta)=\mathrm{Re}\left\{ik\frac{\left(1-e^{i\left(1-\mu\right)k\chi_{*}}\right)}{1-\mu}e^{-ik\eta}\sum\limits_{s}\sigma^{s}(\hat{k},\hat{p})h_{s}^{0}(k)\right\}.$
To lowest order, proper time of the observer equals coordinate time; $\tau=\eta+\mathcal{O}(h)$. Thus, integrating with respect to coordnate time $\eta$ gives the polarization angle relative to the mean angle:
$\Omega^{\left(T\right)}(\eta)=\mathrm{Re}\left\{\left(\sigma^{+}(\hat{k},\hat{p})h_{+}^{0}(k)+\sigma^{\times}(\hat{k},\hat{p})h_{\times}^{0}(k)\right)\frac{\left(e^{ik(1-\mu)\chi_{*}}-1\right)}{\left(1-\mu\right)}e^{-ik\eta}\right\}.$ (70)
In order to calculate $\Omega^{(T)}$, we need to separate the amplitude and phase of each complex factor on the right-hand side of eq. (70). First, the factor that depends on the distance $\chi_{*}$ to the emitter can be written in polar form as
$e^{i\left(1-\mu\right)k\chi_{*}}-1=2ie^{i(1-\mu)k\chi_{*}/2}\sin(\frac{\left(1-\mu\right)}{2}k\chi_{*}).$
Next, the polarization amplitudes $h_{+}^{0}$ and $h_{\times}^{0}$ will in general have different phases. Let $h_{s}^{0}=H_{s}e^{i\alpha_{s}}$, where $H_{s}$ and $\alpha_{s}$ are respectively the real amplitude and phase of polarization component $s$. Let $\mathcal{K}\equiv\{k,\hat{k},H_{+},H_{-},\alpha_{+},\alpha_{-}\}$ denote the unknown state vector defining the tensor mode. $k$ is the angular frequency, $\hat{k}$ the spatial propagation direction, while $H_{\pm}$ and $\alpha_{\pm}$ are the amplitudes and phases, respectively, of the two polarization components. Then, after some algebra, we find that
$\sigma^{+}(\hat{k},\hat{p})h_{+}^{0}(k)+\sigma^{\times}(\hat{k},\hat{p})h_{\times}^{0}(k)=T(\mathcal{K},\hat{p})e^{i\psi(\mathcal{K},\hat{p})},$
with amplitude
$T(\mathcal{K},\hat{p})\equiv{\left({\left(\sigma_{+}H_{+}\right)}^{2}+{\left(\sigma_{\times}H_{\times}\right)}^{2}+2\left(\sigma_{+}H_{+}\right)\left(\sigma_{\times}H_{\times}\right)\cos(\alpha_{+}-\alpha_{\times})\right)}^{1/2}$ (71)
and phase
$\psi(\mathcal{K},\hat{p})\equiv\arctan(\frac{\sigma_{+}H_{+}\sin\alpha_{+}+\sigma_{\times}H_{\times}\ \ \sin\alpha_{\times}}{\sigma_{+}H_{+}\cos\alpha_{+}+\sigma_{\times}H_{\times}\ \ \cos\alpha_{\times}}).$
Finally, the polarization angle takes the sinusoidal form
$\Omega^{\left(T\right)}(\eta)=W(\mathcal{K},\hat{p},\chi_{*})\sin(k\eta+\gamma(\mathcal{K},\hat{p},\chi_{*})),$ (72)
with amplitude
$W(\mathcal{K},\hat{p},\chi_{*})\equiv 2T(\mathcal{K},\hat{p})\frac{\sin(\frac{\left(1-\mu\right)}{2}k\chi_{*})}{1-\mu}$ (73)
and phase
$\gamma(\mathcal{K},\hat{p},\chi_{*})\equiv-\left(\frac{\left(1-\mu\right)}{2}k\chi_{*}+\psi(\mathcal{K},\hat{p})\right).$
The polarization wiggling amplitude $W(\mathcal{K},\hat{p},\chi_{*})$ defines the polarization wiggling response of polarized radiation emitted at distance $\chi_{*}$ from the detector and in direction $\hat{p}$ to a plane gravitational wave with state vector $\mathcal{K}$. The amplitude $W$ and phase $\gamma$ vary with the state $\mathcal{K}$ of the gravitational wave as well as the position of the emitter relative to the detector. We see from eq. (72) that the polarization wiggling frequency is equal to the gravitational wave frequency.
Let us take a look at the conditions under which the amplitude $W$ vanishes. First of all, for inclinations $\mu=\pm 1$, the direction of the emitted radiation is aligned with the direction of the gravitational tensor mode; $\hat{p}=\pm\hat{k}$. From eqs. (68) and (69) follows that $\sigma^{+}(\hat{k},\pm\hat{k})=\sigma^{\times}(\hat{k},\pm\hat{k})=0$, which implies that
$W(\mathcal{K},\pm\hat{k},\chi_{*})=0.$
Hence, polarized radiation is insensitive to gravitational radiation propagating in the same or opposite direction.
Next, there are certain emitter distances that give vanishing $W$ for $|\mu|<1$:
$\chi_{*}=\frac{2\pi N}{\left(1-\mu\right)k},N=1,2,...$
With the exception of these cases as well as other border cases where $T(\mathcal{K},\hat{p})$ vanishes for certain combinations of gravitational wave state $\mathcal{K}$ and polarized radiation direction $\hat{p}$, $W(\mathcal{K},\hat{p},\chi_{*})\neq 0$. Therefore, polarized electromagenetic radiation is in general subject to polarization wiggling in the presence of gravitational radiation.
VIII.4 Wiggle Rate from a Tensor Mode in an Expanding Universe
Let us assume a conformally flat background with ${a(\eta)}^{\prime}\neq 0$ and evaluate the wiggle rate of eq. (64). In order to get analytical results and gain an understanding of the results, we will make the simplifying assumption that the observed electromagnetic radiation is emitted in a matter-dominated spacetime. Although not a realistic model of our universe at later times, we expect that this approximation is a reasonable approximation for tensor modes of smaller scale emitted early enough to be sufficiently damped at the end of the matter-dominate era to not yield significant contributions to the wiggling rate from later times. In addition, it is a useful approximation to get a qualitative understanding of how gravitational tensor modes induce polarization wiggling in an expanding universe.
It is known that a matter-dominated universe has a scale factor
$a(\eta)={\left(\frac{\eta}{\eta_{0}}\right)}^{2},$
where the present conformal time is
$\eta_{0}=\frac{2}{H_{0}\sqrt{\Omega_{m}}},$
and the conformal time of emission is
$\eta_{*}=\eta_{0}\sqrt{a(\eta_{*})}.$
$H_{0}$ is the Hubble constant and $\Omega_{m}$ is the present matter fraction.
This enables us to evaluate the third term on the right-hand side of eq. (64). Reparametrizing the integral in terms of the variable $x\equiv k\eta$, we get
$ik\mu\operatorname*{\int}\limits_{\eta_{*}}^{\eta}d\eta^{\prime}\frac{e^{-ik\left(1-\mu\right)\eta^{\prime}}}{{a(\eta^{\prime})}^{2}}=ik\mu\eta^{2}J,$
where
$J\equiv\operatorname*{\int}\limits_{x_{*}}^{x}dt\frac{e^{-i\left(1-\mu\right)t}}{t^{2}}$
and $x_{*}\equiv k\eta_{*}$. The integral $J$ can be expressed in terms of the exponential integral $E_{2}$ as follows [32]:
$J=\frac{E_{2}(i\left(1-\mu\right)x_{*})}{x_{*}}-\frac{E_{2}(i\left(1-\mu\right)x_{0})}{x_{0}}.$
We may use that $E_{n}(z)=e^{-z}/z+\mathcal{O}(z^{-2})$ for $|z|\gg 1$ [32]. For small-scale modes with $k\eta_{*}\gg 1$, the wiggle rate in the physical frame then becomes
$\tilde{\omega}(\eta)=\mathrm{Re}\left\{ik\left(1+i\mu\right)\left(1-\frac{{e}^{ik(1-\mu)\chi_{*}}}{{a(\eta_{*})}^{2}}\right)e^{-ik\eta}\sum\limits_{s}\sigma^{s}(\hat{k},\hat{p})h_{s}^{0}(k)\right\}$ (74)
to leading order in $1/k{\eta}_{*}$. This can readily be integrated to obtain the polarization angle $\Omega^{(T)}$ as a function of conformal time:
$\Omega^{\left(T\right)}(\eta)=\mathrm{Re}\left\{e^{-ik\eta}(1+i\mu)\left(\frac{{e}^{i\left(1-\mu\right)k\chi_{*}}}{{a(\eta_{*})}^{2}}-1\right)\sum\limits_{s}\sigma^{s}(\hat{k},\hat{p})h_{s}^{0}(k)\right\}.$ (75)
It is evident from eq. (74) that for radiation emitted early, while $a(\eta_{*})\ll 1$, the ${a(\eta_{*})}^{-2}$ terms on the right-hand side of eq. (74) will be dominant, and the wiggle rate depends primarily on the tensor mode amplitude at emission. This confirms our earlier claim that, in an expanding universe, the dominant contributions to the present wiggle rate of polarized radiation emitted at early times come from gravitational fields present at the time of emission. Taking the real part of the complex expression of eq. (75) yields
$\Omega^{\left(T\right)}(\eta)=W(\mathcal{K},\hat{p},\chi_{*})\cos(k\eta+\gamma(\mathcal{K},\hat{p},\chi_{*}))$ (76)
where $T(\mathcal{K},\hat{p})$ is given by eq. (71) above and
$W(\mathcal{K},\hat{p},\chi_{*})\equiv\frac{T(\mathcal{K},\hat{p})\sqrt{1+\mu^{2}}}{{a(\eta_{*})}^{2}}$ (77)
$\gamma(\mathcal{K},\hat{p},\chi_{*})\equiv-\left(\xi+\left(1-\mu\right)k\chi_{*}+\psi(\mathcal{K},\hat{p})\right).$ (78)
We see from eq. (76) that the polarization wigging frequency is equal to the gravitational wave frequency.
VIII.5 Using Polarization Wiggling to Measure the State Parameters of a Plane Gravitational Wave
Given the results of this section, an obvious question is: Can electromagnetic polarization be used to measure all parameters of the state vector $\mathcal{K}$ of a plane gravitational wave?
The state vector $\mathcal{K}\equiv\{k,\hat{k},H_{+},H_{-},\alpha_{+},\alpha_{-}\}$ has 7 independent parameters. Since the frequency of the polarization wiggling induced by a plane gravitational wave is the same as the frequency of the gravitational wave itself, the angular frequency $k$ of a gravitational wave can be measured by simply measuring the frequency of the induced polarization wiggling.
The effect of the other 6 parameters are entangled in the amplitude $W$ and phase $\gamma$. Assuming the polarized radiation is a plane electromagnetic wave, both the wiggling amplitude $W$ and phase $\gamma$ are measurable quantities. It should therefore be sufficient, at least in principle, to determine the remaining 6 parameters of the state vector $\mathcal{K}$ by measuring the wiggling amplitudes and phases of three beams of polarized radiation emitted by emitters at known positions from different directions.
IX Conclusions
Polarization wiggling is a direct gravitational effect on the polarization of electromagnetic radiation. The effect is defined in distinctly operational terms from the perspective of a single observer: In the presence of a gravitational field along the radiation trajectory, it manifests as a rotation of the polarization axis of the observed radiation in the frame of observation between subsequent observation events. Based on general results from an earlier paper [23], we have in this paper studied polarization wiggling induced by weak gravitational fields. Linear gravity was used to analyze the polarization wiggling effect in spacetimes with flat or conformally flat backgrounds.
Spacetimes with conformally flat backgrounds is an extensive class of spacetimes that includes most spacetimes of interest in cosmology. They have one thing in common: their homogeneous backgrounds are conformally related to 4-dimensional Minkowski space. This relationship  gives rise to  a major simplification in the analysis of polarization wiggling: We showed how it is possible to calculate the polarization wiggle rates in the conformal frame and transform the results back to the physical frame. The conformal frame is a much simpler geometry than the physical frame due to its flat Minkowski background, and the wiggle rates in the physical frame can be obtained by simply multiplying the conformal frame results with a scale factor. This simplification arises when electromagnetic radiation is decomposed into plane waves. As electromagnetic plane waves propagate along null geodesics, we can make use of the fact that null geodesics are invariant under conformal transformations and that Maxwell’s equations in a 4-dimensional vacuum are also conformally invariant.
In linear gravity, the gravitational field, represented by the metric perturbation $h_{\mu\nu}$, can be decomposed into scalar, vector and tensor perturbations that are sourced and evolve independently. The polarization wiggle rates induced by scalar, vector and tensor perturbations to the metric are gauge invariant quantities that can be analyzed independently.
Scalar perturbations to the metric do not induce polarization wiggling, so electromagnetic polarization is insensitive to the scalar degrees of freedom of the gravitational field.
Vector perturbations to the metric, on the other hand, may induce polarization wiggling. The polarization axis wiggle rate $\omega^{(V)}$ is a direct measure of the longitudinal gravitomagnetic field $\mathcal{B}$ and thereby the matter currents sourcing it. The contribution to the polarization axis wiggle rate from vector perturbations has two terms. The first contribution is due to the difference in the angular frame dragging rate around the direction of propagation between the frame of emission and the frame of observation. The second contribution comes from the presence of non-stationary gravitomagnetic fields along the radiation trajectory. This implies that, for a stationary gravitomagnetic field, the polarization axis wiggle rate simply measures the difference in angular frame dragging rate between emission and observation. Thus, stationary gravitomagnetic fields will not induce polarization wiggling of any passing electromagnetic radiation if this radiation was emitted and measured at locations with negligible gravitomagnetic fields.
As an example of polarization wiggling induced by vector pertrubations, we analyzed the polarization wiggling effect of a stationary axi-symmetric gravitational field on polarized radiation emitted by an emitter in orbit around the gravitational source. Assuming an emitter with a known orbit around a gravitational source with unknown angular momentum, it was shown that the angular momentum of the gravitational source can be determined by measuring the polarization wiggle rate emitted at three different positions along the orbit.
Tensor perturbations to the metric also induce polarization wiggling. The effect of a single gravitational plane wave (tensor mode) with an arbitrary state of polarization was studied in detail for two cases: A tensor mode propagating on respectively a flat background and in an expanding, matter-dominated universe. In both cases, we showed how the state parameters of the tensor mode can be determined by measurements of the polarization wiggling frequency, amplitude and phase of radiation emitted by emitters at known positions from at least three different directions. In both cases, the polarization axis of the observed radiation will wiggle with a frequency that equals the frequency of the gravitational wave. Both amplitude and phase of the polarization wiggling can be used to determine the properties of the tensor mode. The amplitude and phase of the polarization wiggling vary with the position of the emitter; the polarization wiggle rate is sensitive to both direction of the electromagnetic radiation used for the measurement and the distance to the emitter. The wiggling amplitude vanishes for gravitational waves with wave lengths much larger than the distance to the emitter.
A peculiar feature of the polarization wiggling from a tensor mode in an expanding universe is that polarization wiggling captures the properties of the tensor mode at the time of emission of the electromagnetic radiation. While the tensor mode fades away in the expanding universe, its properties at early times will be preserved in the polarization wiggling of any passing polarized electromagnetic radiation.